OpenStudy (anonymous):

How do I find the greatest common factor of something like 117 and 195? Whats the easiest way?

OpenStudy (anonymous):

write down the prime factors of both GCF is the product of common factors

OpenStudy (anonymous):

List the factors of both numbers 117: 1 3 9 13 39 117 195: 1 3 5 13 15 39 65 195 Biggest one looks to be: 39

OpenStudy (anonymous):

numbers that go into 117?

OpenStudy (anonymous):

117=3*3*13 195=5*3*13 GCF=3*13=39

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

my brain is so slow at this lol

OpenStudy (anonymous):

u got it? slow or fast doesnt matter much

OpenStudy (anonymous):

so I list all the numbers that can go into 117?

OpenStudy (anonymous):

once u learn the technique, with practice u would be fast in calculations too

OpenStudy (anonymous):

I have been out of school for awhile

OpenStudy (anonymous):

I hope one day I can grasp this as well as you!